OpenStudy (anonymous):

find the coordinates of the point C, halfway between the points A(5,1) and B(-2,7)

OpenStudy (mathmate):

The x-coordinate of C is the average of the x-coordinates of A and B. Likewise for the y-coordinate. For example: P(4,6), Q(10,12), M is the mid-point. M((4+10)/2,(6+12)/2)=M(7,9)

OpenStudy (anonymous):

use formula for mid points...that is.. (\[(\frac{ x2+x1 }{ 2 },\frac{ y2+y1 }{ 2 })\]

OpenStudy (anonymous):

so it will be as follows.. \[(\frac{ 5-2 }{ 2 },\frac{ 7+1 }{ 2 })\] \[=(\frac{ 3 }{ 2 },4)\]